OpenStudy (anonymous):

Express in general form- y=-1/2x-9

OpenStudy (anonymous):

add 9 to both sides: y+9= - 1/2 *x ... is it -1/2x or x* (-1/2)

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

ok... if y=(-1/2)*x -9 (i hope this is your problem) add 9 to both sides: y+9=(-1/2)*x multiply both sides by (-2): x=-2*y -18 put all on one side: 2y+x+18=0